OpenStudy (anonymous):

A 6-kg bucket of water is being pulled straight up by a string at a constant speed. At a certain point the speed of the bucket begins to change. The bucket now has an upward constant acceleration of magnitude 3 m/s2. What is the tension in the rope now

OpenStudy (anonymous):

So I know the basic F=ma so 6x9.8=58.8 N, but what about the other part. of 3m/s2?

OpenStudy (anonymous):

i think i understand. 9.8+3=12.8 12.8x6kg=76.8N can someone confirm

OpenStudy (festinger):

Sum of all forces = ma. If you draw a free body diagram you should see that: ma=T-mg. 6(3)=T-6(9.8)

OpenStudy (anonymous):

we have to get the total acceleration to solve the problem. 9.8+3 ok, so i drew the diagram on my paper, worked it differently then what you suggested and go the exact same answer

OpenStudy (anonymous):